# Axial Seamount Eruption Forecasting Experiment

### v2

@@ 1 Introduction
This document describes the Axial Seamount Eruption Forecasting Experiment (EFE), launched within the Geohazards Crises Observatory (www.geohazards-observatory.com) to test and advance physics-based approaches for the prediction of volcanic eruptions. The design of the EFE is inspired by, and emulates, the Financial Bubble Experiment [Sornette2009_FBE, Sornette2010a_FBE, Sornette2010b_FBE, Woodard2010_FBE], which pioneered a transparent, real-world testing framework for forecast validation in the context of financial crises. The motivation for the EFE arises from the limitations of existing methods in operational volcano forecasting. Despite remarkable advances in high-precision, real-time monitoring, there remains no universally accepted or reliable framework for diagnosing and forecasting eruptions as they unfold. A prevailing view within both academic and operational communities is that long-term eruption forecasts are generally not feasible. An important objective in volcanology is therefore to develop physics-based models that provide a sound scientific foundation for eruption forecasting [NASEM2017_VolcanicEruptions].
@@
The Eruption Forecasting Experiment (EFE) is based on a physics‐based failure forecast model that has previously been applied only retrospectively to volcano monitoring datasets [Lei2025_CEE, Lei2025_EPSL]. This experiment represents the first application and testing of the model using real‐time data prior to an eruption, providing an opportunity to objectively evaluate its forecasting performance. Because retrospective analyses are susceptible to various forms of bias—including data snooping, hindsight reinterpretation, and post hoc model adjustment—the EFE is designed as a real‐time forecasting experiment to minimise such biases. Only forecasts that are documented and timestamped in advance of the events can provide a truly unbiased test of predictive capability.
@@
This project applies and extends a suite of physics-based models to diagnose critical transitions in volcanic systems. These methods, developed and tested over the past decade in the broader context of rupture and failure forecasting across diverse geophysical systems [Sornette1995, Saleur1996, Sornette1998, Lei2025_CEE, Lei2025_EPSL, Lei2025_GRL, Lei2025_IJRMMS_Singularity], are here applied to the Axial Seamount volcano. The site provides an exceptional natural laboratory owing to its dense instrumental network [Chadwick2012_NatGeo, Nooner2016_Science, Wilcock2016_Science, Chadwick2022_G3], which includes geodetic measurements of seafloor uplift and continuous seismic monitoring that together capture the deformation and seismic activity. The combination of rich observational datasets and advanced diagnostic tools enables a rigorous, physics-based evaluation of volcanic eruption predictability. The EFE team is continuously developing and validating the forecasting approaches, which will be progressively implemented and enhanced in future releases.
@@ 2 Experimental Design
  • Forecast release schedule. We will issue forecasts on a regular schedule announced in advance on the website of the Geohazards Crisis Observatory (www.geohazards-observatory.com). The first release was on 8 November 2025. The current document represents the second release planned on 22 December 2025. We aim to provide forecast updates once a month, or more frequently as conditions warrant.
  • Continuous monitoring and analysis. We plan to continuously analyse real-time monitoring data from Axial Seamount, including geodetic measurements of seafloor uplift and continuous seismicity records, from the Ocean Observatories Initiative (OOI) Regional Cabled Array (RCA), funded by the US National Science Foundation [Kelley2014_MarineGeology]. Together, these data capture the deformation and seismic activity associated with magmatic processes preceding eruptions.
@@ 3.1 Site description
Axial Seamount is a submarine basaltic volcano located approximately 500 km offshore Oregon, USA [Chadwick2022_G3]. It has experienced three eruptive episodes over the past three decades, in 1998, 2011, and 2015. The volcano has been continuously monitored by a network of seafloor instruments [Nooner2016_Science], and since 2014 by the OOI-RCA, which includes bottom pressure recorders, tiltmeters, and seismometers [Kelley2014_MarineGeology] (see Fig. 1). This unique geophysical dataset provides an unprecedented opportunity to investigate pre-eruptive unrest processes and to test quantitative models for eruption forecasting. The 2015 eruption was previously anticipated about seven months in advance, within a one‐year window, based on pattern recognition in the geodetic time series [Nooner2016_Science]. This forecast was formulated empirically from the observed inflation pattern rather than derived from a physics‐based model. However, subsequent applications of this technique have not produced reliable forecasts due to variable inflation rates [Chadwick2024_AGU], highlighting the need for physics‐based forecasting frameworks [Acocella2024_NatRevEarthEnv]. Within the current EFE project, we aim to perform ex-ante prospective forecasting of the next eruption of this volcano using our new physics-based failure forecast model with the continuous monitoring datasets of seafloor uplift and seismicity. All datasets in this version include observations updated to 19 December 2025.
@@ 3.2 Seafloor uplift data
Seafloor uplift measurements at Axial Seamount are obtained in real time from a network of four high-precision bottom pressure recorders (BPRs) deployed in and around the caldera [Chadwick2022_G3]. The cabled instruments record absolute pressure at a rate of 1 Hz. The raw pressure data are converted to depth and then de-tided using harmonic tide models. The drift of the OOI-RCA pressure sensors is constrained to less than 0.5 cm yr-1, based on comparisons with self-calibrating BPR instruments and remotely operated vehicle-based relative pressure measurements referenced to a stable site [Sasagawa2016_ESS, Sasagawa2021_AGU, Chadwick2022_G3, Sullivan2025_AGU]. To suppress oceanographic noise (with amplitudes of about $\pm$5 cm), differential BPR records are also computed by subtracting measurements at a reference station located outside the caldera (MJ03E in Fig. 1) from the station at the center of the caldera (MJ03F in Fig. 1), reducing the residual noise to about $\pm$1 cm and better isolating the true geophysical signal of volcanic inflation and deflation [Chadwick2022_G3]. It is important to note, however, that the differential BPR data is relative (MJ03F relative to MJ03E), and it likely includes some (small) component of slip across the seismically active eastern caldera fault. Fig. 2 shows the temporal evolution of seafloor displacement inferred from the single-station BPR record at the centre of the caldera and from differential BPR measurements.
Figure 2: Seafloor uplift measurements at Axial Seamount were obtained from the OOI–RCA bottom pressure recorder (BPR) data repository [Nooner2016_Science, Chadwick2022_G3]. (a) Time series recorded by the bottom pressure recorder (BPR) at the caldera centre (MJ03F in Fig. 1). (b) Differential BPR data derived by subtracting measurements at the reference site east of the caldera (MJ03E in Fig. 1) from those at the caldera centre (MJ03F). The major offset on the left side of the plots is the deflation that occurred during the 2015 eruption.
@@ 3.3 Seismicity data
Seismic data are recorded by the OOI-RCA at Axial Seamount [Wilcock2016_Science]. Earthquake detection is performed automatically using a short-term/long-term root-mean-square trigger algorithm applied to traces high-pass filtered at 4 Hz. Triggered signals from at least eight channels across four stations are grouped into a single event within a 1.5 s window to ensure robust multi-station detections. Automated classifiers exclude fin-whale calls and (co-eruption) seafloor explosions based on spectral ratios and waveform envelope correlations. P- and S-phase arrivals are refined using polarisation criteria and a kurtosis-based picker, and events are initially located with HYPOINVERSE (Klein, 2022) using a one-dimensional velocity model derived from caldera tomography. The formal absolute location uncertainties are typically $\sim$0.2 km horizontally and $\sim$0.3 km vertically.
For each event, P- and S-wave seismic moments $M_{0}$ were estimated from phase amplitudes following established techniques for marine microearthquakes, and the moment magnitude $M_{\mathrm{w}}$ was computed as [Hanks1979_JGR]:
@@
Figure 3: Seismic measurements at Axial Seamount from the OOI-RCA Earthquake Catalogue [Wilcock2016_Science, Wilcock2017_IEDA]. (a) Magnitude-time sequence of earthquakes at Axial Seamount; the red line indicates the completeness magnitude of the catalogue. (b) Time series of cumulative Benioff strain derived from equation (2) with $q=1/2$.
@@
where the energy release $E$ of the $i$th earthquake is estimated from $\log_{10}E=1.5\,M_{\mathrm{w}}+4.8$ [Kanamori1977_JGR] and $N(t)$ is the cumulative number of earthquakes up to time $t$. The power exponent $q$ controls the relative weighting of small versus large events: $q=0$ corresponds to the cumulative event count, $q=1/3$ yields a measure proportional to the fault length, $q=1/2$ gives a quantity scaling with the square root of seismic energy, $q=2/3$ relates to the rupture area, and $q=1$ leads to the cumulative energy release. Here, we adopt $q=1/2$, corresponding to the so-called Benioff strain [Bufe1993_JGR, Sornette1995], with the resulting time series shown in Fig. 3b. Data gaps in the seismic catalogue were filled using piecewise cubic Hermite interpolation, which preserves the local trends without introducing artificial oscillations; tests with linear interpolation yielded similar results.
@@ 4.1 Theory
Our theory for volcano forecasting builds on the log-periodic power-law singularity (LPPLS) formulation, which has proven effective in describing rupture dynamics in heterogeneous geophysical systems and has recently been validated, in retrospect, using a global dataset of 109 geohazard events [Lei2025_CEE, Lei2025_EPSL, Lei2025_GRL, Lei2025_IJRMMS_Singularity]. The LPPLS model captures two essential ingredients of rupture dynamics: (i) the positive feedback mechanism, represented by a general power law acceleration that leads to a super-exponential trajectory culminating in a finite-time singularity, and (ii) the partial breakdown of continuous scale invariance into discrete scale invariance, which arises from the intermittent damage and rupture processes in heterogeneous materials. Importantly, the LPPLS model leverages the oscillatory patterns inherent in rupture dynamics, transforming what is traditionally perceived as noise into valuable predictive information. The first-order expansion of the general LPPLS formulation is given by [Sornette1995, Lei2025_CEE]:
@@ 4.2 Procedures
The LPPLS model involves seven parameters, $\boldsymbol{\theta}=\{t_{c},\,m,\,\omega,\,A,\,B,\,C,\,\phi\}$, with the critical time $t_{c}$ being the primary parameter of interest. A stable and robust calibration scheme [Filimonov2013_PhysicaA] is employed to fit the LPPLS model to the monitoring datasets of Axial Seamount in order to infer $t_{c}$. For each monitoring dataset, the time series is extracted over the interval $t_{0},\,t_{2}]$, where $t_{0}$ is set to 19 May 2015, marking the end of the during-eruption period identified from deformation measurements [Nooner2016_Science] and earthquake focal mechanism analyses [Levy2018_Geology], and $t_{2}$ corresponds to the current time when the forecast is made. The optimal starting time $t_{1}$ ($t_{0}\leq t_{1}<t_{2}$) for LPPLS calibration is then determined using the Lagrange regularisation approach [Demos2019_PhysicaA], providing an estimate of the potential onset of the crisis (around 2019–2020; specific dates will be reported in the forecast results). The LPPLS model is subsequently calibrated using the data within $t_{1},\,t_{2}]$ to estimate $t_{c}$ through a nonlinear fitting procedure [Filimonov2013_PhysicaA].
@@
  • Bootstrap. Bootstrapping is performed by resampling the residuals from the ordinary least-squares LPPLS fit to generate multiple synthetic datasets, each of which is used to recalibrate the model. The resulting ensemble of $t_{c}$ estimates defines the predictive distribution, which is obtained by applying an adaptive kernel density estimation to the ensemble [Cranmer2001_KernelEstimation].
  • Quantile regression. Quantile regression is used to estimate the variability of the LPPLS fit across multiple quantile levels [Zhang2016_PLOSOne], capturing a range of possible trajectories consistent with the observation data. The ensemble of these quantile-based estimates forms another predictive distribution of $t_{c}$, constructed through an adaptive kernel density estimation [Cranmer2001_KernelEstimation].
  • Modified profile likelihood. The modified profile likelihood method constructs a likelihood function for $t_{c}$ while integrating the effects of the other parameters. It provides a statistically rigorous way to estimate $t_{c}$ and its confidence interval by explicitly accounting for the uncertainty of other parameters in the LPPLS model [Filimonov2017_QuantFin]. The resulting likelihood function forms the basis for constructing the predictive distribution of $t_{c}$.
@@
Although these three approaches differ in implementation, they are all rooted in the same LPPLS theoretical framework and provide complementary measures of forecast uncertainty. Bootstrapping captures sampling uncertainty by generating many alternative realisations of the data through residual resampling. Quantile regression probes different parts of the residual distribution to reflect asymmetric or heterogeneous fluctuations [Zhang2016_PLOSOne]. The modified profile likelihood focuses directly on how tightly the calibration can constrain the critical time $t_{c}$, propagating parametric interactions and structural uncertainty into a likelihood-based confidence measure [Filimonov2017_QuantFin]. Together, these methods illuminate distinct facets of uncertainty and yield a more complete characterisation of the predictive distribution of $t_{c}$.
@@ Acknowledgements
This material is based upon work supported by the Ocean Observatories Initiative (OOI), a major facility fully funded by the U.S. National Science Foundation under Cooperative Agreement No. 2244833, and the Woods Hole Oceanographic Institution OOI Program Office. Q.L. and D.S. acknowledge partial support from the Norwegian Water Resources and Energy Directorate (NVE). D.S. acknowledges funding by the National Natural Science Foundation of China (Grant No. T2350710802 and U2039202), and the Center for Computational Science and Engineering at Southern University of Science and Technology. W.W.C. and S.N. acknowledge funding from US National Science Foundation awards (Grant No. OCE-2226488 and 2226445).

### v3

@@ 1 Introduction
This document describes the Axial Seamount Eruption Forecasting Experiment (EFE), launched within the Geohazards Crises Observatory (www.geohazards-observatory.com) to test and advance physics-based approaches for the prediction of volcanic eruptions. The design of the EFE is inspired by, and emulates, the Financial Bubble Experiment [26, 24, 25, 32], which pioneered a transparent, real-world testing framework for forecast validation in the context of financial crises. The motivation for the EFE arises from the limitations of existing methods in operational volcano forecasting. Despite remarkable advances in high-precision, real-time monitoring, there remains no universally accepted or reliable framework for diagnosing and forecasting eruptions as they unfold. A prevailing view within both academic and operational communities is that long-term eruption forecasts are generally not feasible. An important objective in volcanology is therefore to develop physics-based models that provide a sound scientific foundation for eruption forecasting [18].
@@
The Eruption Forecasting Experiment (EFE) is based on a physics‐based failure forecast model that has previously been applied only retrospectively to volcano monitoring datasets [16, 14]. This experiment represents the first application and testing of the model using real‐time data prior to an eruption, providing an opportunity to objectively evaluate its forecasting performance. Because retrospective analyses are susceptible to various forms of bias—including data snooping, hindsight reinterpretation, and post hoc model adjustment—the EFE is designed as a real‐time forecasting experiment to minimise such biases. Only forecasts that are documented and timestamped in advance of the events can provide a truly unbiased test of predictive capability.
@@
This project applies and extends a suite of physics-based models to diagnose critical transitions in volcanic systems. These methods, developed and tested over the past decade in the broader context of rupture and failure forecasting across diverse geophysical systems [23, 20, 27, 16, 14, 13, 15], are here applied to the Axial Seamount volcano. The site provides an exceptional natural laboratory owing to its dense instrumental network [3, 19, 30, 4], which includes geodetic measurements of seafloor uplift and continuous seismic monitoring that together capture the deformation and seismic activity. The combination of rich observational datasets and advanced diagnostic tools enables a rigorous, physics-based evaluation of volcanic eruption predictability. The EFE team is continuously developing and validating the forecasting approaches, which will be progressively implemented and enhanced in future releases.
@@ 2 Experimental Design
  • Forecast release schedule. We will issue forecasts on a regular schedule announced in advance on the website of the Geohazards Crisis Observatory (www.geohazards-observatory.com). The first release was issued on 8 November 2025 and the second on 22 December 2025. This document constitutes the third release, planned for 30 January 2026. We aim to provide forecast updates once a month, or more frequently as conditions warrant.
  • Continuous monitoring and analysis. We plan to continuously analyse real-time monitoring data from Axial Seamount, including geodetic measurements of seafloor uplift and continuous seismicity records, from the Ocean Observatories Initiative (OOI) Regional Cabled Array (RCA), funded by the US National Science Foundation [12]. Together, these data capture the deformation and seismic activity associated with magmatic processes preceding eruptions.
@@ 3.1 Site description
Axial Seamount is a submarine basaltic volcano located approximately 500 km offshore Oregon, USA [4]. It has experienced three eruptive episodes over the past three decades, in 1998, 2011, and 2015. The volcano has been continuously monitored by a network of seafloor instruments [19], and since 2014 by the OOI-RCA, which includes bottom pressure recorders, tiltmeters, and seismometers [12] (see Fig. 1). This unique geophysical dataset provides an unprecedented opportunity to investigate pre-eruptive unrest processes and to test quantitative models for eruption forecasting. The 2015 eruption was previously anticipated about seven months in advance, within a one‐year window, based on pattern recognition in the geodetic time series [19]. This forecast was formulated empirically from the observed inflation pattern rather than derived from a physics‐based model. However, subsequent applications of this technique have not produced reliable forecasts due to variable inflation rates [5], highlighting the need for physics‐based forecasting frameworks [1]. Within the current EFE project, we aim to perform ex-ante prospective forecasting of the next eruption of this volcano using our new physics-based failure forecast model with the continuous monitoring datasets of seafloor uplift and seismicity. All datasets in this version include observations updated to 24 January 2026.
@@ 3.2 Seafloor uplift data
Seafloor uplift measurements at Axial Seamount are obtained in real time from a network of four high-precision bottom pressure recorders (BPRs) deployed in and around the caldera [4]. The cabled instruments record absolute pressure at a rate of 1 Hz. The raw pressure data are converted to depth and then de-tided using harmonic tide models. The drift of the OOI-RCA pressure sensors is constrained to less than 0.5 cm yr-1, based on comparisons with self-calibrating BPR instruments and remotely operated vehicle-based relative pressure measurements referenced to a stable site [21, 22, 4, 28]. To suppress oceanographic noise (with amplitudes of about $\pm$5 cm), differential BPR records are also computed by subtracting measurements at a reference station located outside the caldera (MJ03E in Fig. 1) from the station at the center of the caldera (MJ03F in Fig. 1), reducing the residual noise to about $\pm$1 cm and better isolating the true geophysical signal of volcanic inflation and deflation [4]. It is important to note, however, that the differential BPR data is relative (MJ03F relative to MJ03E), and it likely includes some (small) component of slip across the seismically active eastern caldera fault. Fig. 2 shows the temporal evolution of seafloor displacement inferred from the single-station BPR record at the centre of the caldera and from differential BPR measurements.
Figure 2: Seafloor uplift measurements at Axial Seamount were obtained from the OOI–RCA bottom pressure recorder (BPR) data repository [19, 4] (updated as of 24 January 2026). (a) Time series recorded by the bottom pressure recorder (BPR) at the caldera centre (MJ03F in Fig. 1). (b) Differential BPR data derived by subtracting measurements at the reference site east of the caldera (MJ03E in Fig. 1) from those at the caldera centre (MJ03F). The major offset on the left side of the plots is the deflation that occurred during the 2015 eruption.
@@ 3.3 Seismicity data
Seismic data are recorded by the OOI-RCA at Axial Seamount [30]. Earthquake detection is performed automatically using a short-term/long-term root-mean-square trigger algorithm applied to traces high-pass filtered at 4 Hz. Triggered signals from at least eight channels across four stations are grouped into a single event within a 1.5 s window to ensure robust multi-station detections. Automated classifiers exclude fin-whale calls and (co-eruption) seafloor explosions based on spectral ratios and waveform envelope correlations. P- and S-phase arrivals are refined using polarisation criteria and a kurtosis-based picker, and events are initially located with HYPOINVERSE (Klein, 2022) using a one-dimensional velocity model derived from caldera tomography. The formal absolute location uncertainties are typically $\sim$0.2 km horizontally and $\sim$0.3 km vertically.
For each event, P- and S-wave seismic moments $M_{0}$ were estimated from phase amplitudes following established techniques for marine microearthquakes, and the moment magnitude $M_{\mathrm{w}}$ was computed as [10]:
@@
Figure 3: Seismic measurements at Axial Seamount from the OOI-RCA Earthquake Catalogue [30, 31] (updated as of 24 January 2026). (a) Magnitude-time sequence of earthquakes at Axial Seamount; the red line indicates the completeness magnitude of the catalogue. (b) Time series of cumulative Benioff strain derived from equation (2) with $q=1/2$.
@@
where the energy release $E$ of the $i$th earthquake is estimated from $\log_{10}E=1.5\,M_{\mathrm{w}}+4.8$ [11] and $N(t)$ is the cumulative number of earthquakes up to time $t$. The power exponent $q$ controls the relative weighting of small versus large events: $q=0$ corresponds to the cumulative event count, $q=1/3$ yields a measure proportional to the fault length, $q=1/2$ gives a quantity scaling with the square root of seismic energy, $q=2/3$ relates to the rupture area, and $q=1$ leads to the cumulative energy release. Here, we adopt $q=1/2$, corresponding to the so-called Benioff strain [2, 23], with the resulting time series shown in Fig. 3b. Data gaps in the seismic catalogue were filled using piecewise cubic Hermite interpolation, which preserves the local trends without introducing artificial oscillations; tests with linear interpolation yielded similar results.
@@ 4.1 Theory
Our theory for volcano forecasting builds on the log-periodic power-law singularity (LPPLS) formulation, which has proven effective in describing rupture dynamics in heterogeneous geophysical systems and has recently been validated, in retrospect, using a global dataset of 109 geohazard events [16, 14, 13, 15]. The LPPLS model captures two essential ingredients of rupture dynamics: (i) the positive feedback mechanism, represented by a general power law acceleration that leads to a super-exponential trajectory culminating in a finite-time singularity, and (ii) the partial breakdown of continuous scale invariance into discrete scale invariance, which arises from the intermittent damage and rupture processes in heterogeneous materials. Importantly, the LPPLS model leverages the oscillatory patterns inherent in rupture dynamics, transforming what is traditionally perceived as noise into valuable predictive information. The first-order expansion of the general LPPLS formulation is given by [23, 16]:
@@ 4.2 Procedures
The LPPLS model involves seven parameters, $\boldsymbol{\theta}=\{t_{c},\,m,\,\omega,\,A,\,B,\,C,\,\phi\}$, with the critical time $t_{c}$ being the primary parameter of interest. A stable and robust calibration scheme [9] is employed to fit the LPPLS model to the monitoring datasets of Axial Seamount in order to infer $t_{c}$. For each monitoring dataset, the time series is extracted over the interval $t_{0},\,t_{2}]$, where $t_{0}$ is set to 19 May 2015, marking the end of the during-eruption period identified from deformation measurements [19] and earthquake focal mechanism analyses [17], and $t_{2}$ corresponds to the current time when the forecast is made. The optimal starting time $t_{1}$ ($t_{0}\leq t_{1}<t_{2}$) for LPPLS calibration is then determined using the Lagrange regularisation approach [7], providing an estimate of the potential onset of the crisis (around 2019–2020; specific dates will be reported in the forecast results). The LPPLS model is subsequently calibrated using the data within $t_{1},\,t_{2}]$ to estimate $t_{c}$ through a nonlinear fitting procedure [9].
@@
  • Bootstrap. Bootstrapping is performed by resampling the residuals from the ordinary least-squares LPPLS fit to generate multiple synthetic datasets, each of which is used to recalibrate the model. The resulting ensemble of $t_{c}$ estimates defines the predictive distribution, which is obtained by applying an adaptive kernel density estimation to the ensemble [6].
  • Quantile regression. Quantile regression is used to estimate the variability of the LPPLS fit across multiple quantile levels [33], capturing a range of possible trajectories consistent with the observation data. The ensemble of these quantile-based estimates forms another predictive distribution of $t_{c}$, constructed through an adaptive kernel density estimation [6].
  • Modified profile likelihood. The modified profile likelihood method constructs a likelihood function for $t_{c}$ while integrating the effects of the other parameters. It provides a statistically rigorous way to estimate $t_{c}$ and its confidence interval by explicitly accounting for the uncertainty of other parameters in the LPPLS model [8]. The resulting likelihood function forms the basis for constructing the predictive distribution of $t_{c}$.
@@
Although these three approaches differ in implementation, they are all rooted in the same LPPLS theoretical framework and provide complementary measures of forecast uncertainty. Bootstrapping captures sampling uncertainty by generating many alternative realisations of the data through residual resampling. Quantile regression probes different parts of the residual distribution to reflect asymmetric or heterogeneous fluctuations [33]. The modified profile likelihood focuses directly on how tightly the calibration can constrain the critical time $t_{c}$, propagating parametric interactions and structural uncertainty into a likelihood-based confidence measure [8]. Together, these methods illuminate distinct facets of uncertainty and yield a more complete characterisation of the predictive distribution of $t_{c}$.
@@ Acknowledgements
This material is based upon work supported by the Ocean Observatories Initiative (OOI), a major facility fully funded by the U.S. National Science Foundation under Cooperative Agreement No. 2244833, and the Woods Hole Oceanographic Institution OOI Program Office. Q.L. is grateful for support from the European Research Council (ERC) under the European Union’s Horizon Europe programme (ERC Consolidator Grant, grant no. 101232311) for the project “Unified framework for modelling progressive to catastrophic failure in fractured media (FORECAST)”. Q.L. and D.S. acknowledge partial support from the Norwegian Water Resources and Energy Directorate (NVE). D.S. acknowledges funding by the National Natural Science Foundation of China (Grant No. T2350710802 and U2039202), and the Center for Computational Science and Engineering at Southern University of Science and Technology. W.W.C. and S.N. acknowledge funding from US National Science Foundation awards (Grant No. OCE-2226488 and 2226445).